# On the Convergence of Loss and Uncertainty-based Active Learning Algorithms

### v3

@@ 1 Introduction
In practice, when training machine learning models for prediction tasks (classification or regression), one often has access to an abundance of unlabeled data, while obtaining the corresponding labels may entail high costs. This may especially be the case in fields like computer vision, natural language processing, and speech recognition.
Active learning algorithms are designed to efficiently learn a prediction model by employing a label acquisition, with the goal of minimizing the number of labels used to train an accurate prediction model.
@@
There is a large body of work on convergence of SGD algorithms, e.g. see Bubeck (2015); Nesterov (2018). These results are established for SGD algorithms under either constant, diminishing or adaptive step sizes. Recently, Loizou et al. (2021), studied SGD with the stochastic Polyak’s step size, depending on the ratio of the loss and the squared gradient of the loss of a point. Our work is different as we consider convergence of SGD under sampling of points.
@@ 2 Problem Statement
We consider the projected SGD algorithm defined as: given an initial value $\theta_{1}\in\Theta$,
@@
where $\ell:\mathcal{X}\times\mathcal{Y}\times\Theta\rightarrow\mathbb{R}$ is a training loss function, $z_{t}$ is a stochastic step size with mean $\zeta(x_{t},y_{t},\theta_{t})$ for some function $\zeta:\mathcal{X}\times\mathcal{Y}\times\Theta\mapsto\mathbb{R}_{+}$, $\Theta_{0}\subseteq\Theta$, and $\mathcal{P}_{\Theta_{0}}$ is the projection function, i.e., $\mathcal{P}_{\Theta_{0}}(u)=\arg\min_{v\in\Theta_{0}}||u-v||$. Unless specified otherwise, we consider the case $\Theta_{0}=\Theta$, which requires no projection. For binary classification tasks, we assume $\mathcal{Y}=\{-,1,1\}$. For every $t>0$, we define $\bar{\theta}_{t}=(1/t)\sum_{s=1}^{t}\theta_{s}$.
@@ 3.1 Constant-Weight Sampling
$\mathbb{E}\left[\ell(yx^{\top}\bar{\theta}_{n})\right]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(y_{t}x_{t}^{\top}\theta_{t})\right]\leq\frac{R^{2}S^{2}}{\beta}\frac{1}{n}.$
@@ 4 Numerical Results
In active learning applications, the true loss of a point cannot be computed before the corresponding label is obtained. Hence, in practice we do not know the true loss at the moment of making the sampling decision. Therefore, we assess the effect of using a loss estimator, instead of using the true loss values. We use a Random Forest regressor to estimate absolute error loss based on the same set of features as the target model and the target’s model prediction as an extra feature. We retrain this estimator on every sampling step using the labeled points observed so far. Figure 2 demonstrates that AWS-PA with the estimated absolute error losses performs similarly on all datasets to AWS-PA with the true absolute error losses. Moreover, for a majority of the datasets, the two variants of AWS-PA achieve similar data sampling rates; this is shown Appendix B.3 along with further discussion.
@@ 5 Conclusion
In this paper, we have provided convergence rate guarantees for both loss and uncertainty-based active learning algorithms under various assumptions. Furthermore, we introduced the novel Adaptive-Weight Sampling (AWS) algorithm that combines sampling with an adaptive size, conforming to stochastic Polyak’s step size in expectation, and demonstrated its convergence rate guarantee, contingent on a condition related to the sampling probability function. For future research, it would be interesting to apply the methodologies presented in this paper to obtain additional convergence rate results. Additionally, exploring a theoretical study on the impact of bias and noise in the loss estimator, used for evaluating the sampling probability function, on the convergence properties of algorithms could open up a valuable avenue for investigation.

### v4

@@ 1 Introduction
In practice, when training machine learning models for prediction tasks (classification or regression), one often has access to an abundance of unlabeled data, while obtaining the corresponding labels may entail high costs. This may especially be the case in fields like computer vision, natural language processing, and speech recognition. Active learning algorithms are designed to efficiently learn a prediction model by employing a label acquisition, with the goal of minimizing the number of labels used to train an accurate prediction model.
@@
There is a large body of work on convergence of SGD algorithms, e.g. see Bubeck (2015) and Nesterov (2018). These results are established for SGD algorithms under either constant, diminishing or adaptive step sizes. Recently, Loizou et al. (2021), studied SGD with the stochastic Polyak’s step size, depending on the ratio of the loss and the squared gradient of the loss of a point. Our work proposes an adaptive-window sampling algorithm and provides its convergence analysis, with the algorithm defined as SGD with a sampling of points and an adaptive step size update that conform to the stochastic Polyak’s step size in expectation. This is unlike to the adaptive step size SGD algorithm by Loizou et al. (2021) which does not use sampling.
@@ 2 Problem Statement
We consider the setting of streaming algorithms where a machine learning model parameter $\theta_{t}$ is updated sequentially, upon encountering each data point, with $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\in\mathcal{X}\times\mathcal{Y}$ denoting the sequence of data points with the corresponding labels, assumed to be independent and identically distributed with distribution $\mathcal{D}$. Specifically, we consider the class of projected SGD algorithms defined as: given an initial value $\theta_{1}\in\Theta$,
@@
where $\ell:\mathcal{X}\times\mathcal{Y}\times\Theta\rightarrow\mathbb{R}$ is a training loss function, $z_{t}$ is a stochastic step size with mean $\zeta(x_{t},y_{t},\theta_{t})$ for some function $\zeta:\mathcal{X}\times\mathcal{Y}\times\Theta\mapsto\mathbb{R}_{+}$, $\Theta_{0}\subseteq\Theta$, and $\mathcal{P}_{\Theta_{0}}$ is the projection function, i.e., $\mathcal{P}_{\Theta_{0}}(u)=\arg\min_{v\in\Theta_{0}}||u-v||$. Unless specified otherwise, we consider the case $\Theta_{0}=\Theta$, which requires no projection. For binary classification tasks, we assume $\mathcal{Y}=\{-1,1\}$. For every $t>0$, we define $\bar{\theta}_{t}=(1/t)\sum_{s=1}^{t}\theta_{s}$.
By defining the distribution of the stochastic step size $z_{t}$ in Equation (1) appropriately, we can accommodate different active learning and data subset selection algorithms. In the context of active learning algorithms, at each step $t$, the algorithm observes the value of $x_{t}$ and decides whether or not to observe the value of the label $y_{t}$. The value of $z_{t}$ determine whether or not we observe the label $y_{t}$. Deciding not to observe the value of the label $y_{t}$ implies the step size $z_{t}$ of value zero (not updating the machine learning model).
@@ 3.1 Constant-Weight Sampling
$\mathbb{E}\left[\ell(yx^{\top}\bar{\theta}_{n})\right]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(y_{t}x_{t}^{\top}\theta_{t})\right]\leq\frac{R^{2}S^{2}}{\beta}\frac{1}{n},$
where $(x,y)$ is an independent sample of a labeled data point from $\mathcal{D}$.
@@ 4 Numerical Results
The implementation of AWS-PA algorithm along with all the other code run the experimental setup that is described in this section is available at https://www.github.com/facebookresearch/AdaptiveWeightSampling.
@@
In active learning applications, the true loss of a point cannot be computed before the corresponding label is obtained. Hence, in practice we do not know the true loss at the moment of making the sampling decision. Therefore, we assess the effect of using a loss estimator, instead of using the true loss values. We use a Random Forest regressor to estimate absolute error loss based on the same set of features as the target model and the target’s model prediction as an extra feature. We retrain this estimator on every sampling step using the labeled points observed so far.
Figure 2 demonstrates that AWS-PA with the estimated absolute error losses performs similarly on all datasets to AWS-PA with the true absolute error losses. Moreover, for a majority of the datasets, the two variants of AWS-PA achieve similar data sampling rates; this is shown Appendix B.3 along with further discussion.
@@ 5 Conclusion
We have provided convergence rate guarantees for loss and uncertainty-based active learning algorithms under various assumptions. Furthermore, we introduced the novel Adaptive-Weight Sampling (AWS) algorithm that combines sampling with an adaptive size, conforming to stochastic Polyak’s step size in expectation, and demonstrated its convergence rate guarantee, contingent on a condition related to the sampling probability function.
For future research, it would be interesting to establish tight convergence rates for the training loss function and the sampling cost, especially comparing policies using sampling with a constant probability with those using adaptive loss-based sampling probabilities. It would be interesting to explore adaptive-weight sampling algorithms with adaptive sizes different than those studied in this paper. Additionally, exploring a theoretical study on the impact of bias and noise in the loss estimator, used for evaluating the sampling probability function, on the convergence properties of algorithms could open up a valuable avenue for investigation.
@@ B.3 Further Details on the Robustness of AWS-PA to Loss Estimation
Alternative absloss estimators
In Figure 2 we shared results of AWS-PA using a Random Forest regressor to estimate the absolute loss. Figure 9 shows the results the credit dataset of an otherwise identical experimental setup where we have replaced the Random Forest regressor absloss estimator with an MLP neural network.
Figure 9: Credit dataset with an MLP neural network loss estimator.